a) Bid Price of New Zealand Dollar - JP Morgan Bank USD0.6533 and Well Fargo USD0.6503

Ask Price of New Zealand Dollar - JP Morgan Bank USD0.6563 and Well Fargo USD0.6523

Justify whether locational arbitrage is possible. If so, explain the steps involved in locational arbitrage, and estimate the profit from this arbitrage if you had USD1,000,000 to use. Discuss market forces factors that would occur to eliminate any further possibilities of locational arbitrage. (6 marks)

b) Currency Pair Quoted Price Value of Canadian Dollar (CAD) in US Dollars (USD) USD0.7381

Value of New Zealand Dollar (NZD) in US Dollars (USD) USD0.6537

Value of Canadian Dollar in New Zealand Dollars (NZD) NZD1.1305

Referring to the information given, explain the steps that would reflect triangular arbitrage. If you had USD2,500,000 to use in triangular arbitrage, compute the profit from this strategy. Justify the steps taken to prevent any further possibilities of triangular arbitrage. (7 marks)

c) Given the following information, calculate the yield (percentage return) to a U.S. investor who used covered interest arbitrage. Assume the investor invests USD2,000,000 determine the profit or loss in this transaction. You are required to analyse the market forces determinants occur in order to protect from any further possibilities of covered interest arbitrage.

Currency pairs and Interest Rate Quoted Price Spot Rate of Canadian Dollar (CAD) USD0.7381

90-day forward rate of Canadian dollar USD0.7252

90-day Canadian interest rate 3.25%

90-day U.S. interest rate 1.75%

d) Assume that interest rate parity exists. You expect that the one-year nominal interest rate in the U.S. is 1.995%, while the one-year nominal interest rate in Australia is 3.695%. The spot rate of the Australian dollar is USD0.6939. You will need 15 million Australian Dollars in one year. Today, you purchase a one-year forward contract in Australian Dollars. Estimate how many U.S. Dollars (USD) will you need in one year to fulfill your forward contract.
Answer #1

a)

In JP Morgan Bank, $/NZ $ = Bid 0.6533-Ask 0.6563

In Well Fargo, $/NZ $ = Bid 0.6503-Ask 0.6523

We can Buy NZ $ at Ask Rate and Sell NZ $ at Bid Rate

Therefore, there are 2 alternatives:

  1. Buying from JP Morgan @ $0.6563 and Selling to Well Fargo @ $0.6503
  2. Buying from Well Fargo @ $0.6523 and Selling to JP Morgan @ $0.6533

Out of both, (i) will result in Loss and (ii) will result in Profit

Therefore, Steps for Arbitrage:

  1. Buy NZ $ from Well Fargo @ $0.6523. For $1000000, 1000000/0.6523 = NZ $1533036.95 will be received
  2. Sell NZ $ to JP Morgan @ $0.6533. For NZ $1533036.95, 1533036.95*0.6533 = $1001533.04 will be received

Therefore, Arbitrage Profit = 1001533.04-1000000 = $1533.04

Note: As per GUIDELINES, we are supposed to Answer ONLY 1 QUESTION.